OpenStudy (anonymous):

What is 2pi / (pi /2)

OpenStudy (anonymous):

simplified

OpenStudy (anonymous):

@CliffSedge

OpenStudy (anonymous):

how many times does half of something go into 2 of something?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

\[\frac{ 2\pi }{ 1 }*\frac{ 2 }{ \pi }\] then cancel out the Pies!

OpenStudy (anonymous):

so its 4?

OpenStudy (anonymous):

\[\frac{ 2 }{ 1 }*\frac{ 2 }{ 1 } = 4\] yep

OpenStudy (anonymous):

thnx